There are two stable isotopes of calcium: calcium -40(39.96amu) and calcium -46(45.95amu). Using the average atomic mass from the periodic table, determine which isotope has a higher percent abundance.
Answer #1

Avarage atomic mass of Ca = 40.078amu

(mass fraction of Ca-40 × atomic mass of Ca-40) + (mass fraction of Ca-46 × atomic mass of Ca- 40) = Avarage atomic mass of Ca

mass fraction of Ca-40 is x

mass fraction of Ca-46 is 1-x

( x × 39.96 amu ) + ( (1-x) × 45.95amu) = 40.078amu

x39.96amu + 45.95amu - x 45.95amu = 40.078amu

- x 5.99 = - 5.872amu

x = 0.9803

1-x = 1- 0.9803 = 0.0197

percent abundance of Calcium-40 = 98.03%

percent abundance of calcium-46 = 1.97%

Therefore,

Calcium - 40 has a higher percent abundance.